OpenStudy (anonymous):

2x - y - 4 = 0 3x + y - 9 = 0 What is the solution set of the given system?

OpenStudy (anonymous):

if you add this equations, you get just simple equations. 1) you must find x from simple eq. 2) you must find y using known x, by using one of this eq.

OpenStudy (blurbendy):

5x - 13 = 0 5x = 13 x = 13 /5 2(13/5) - y - 4 = 0 5.2 - y - 4 = 0 1.2 -y = 0 -y = -1.2 y = 1.2 y = 6/5

OpenStudy (blurbendy):

Just add the first equation to the second equation to get to the first line in my answer.